OpenStudy (gudetamas):

WILLING TO LEARN, PLEASE HELP Which statement about the quadratic functions below is false? f(x)= -3/4x^2+6 g(x)= -2x^2 - 5 h(x) = 1/4x^2+1 a. the graphs of two of these functions have a minimum point b.the graphs of all these functions have the same axis of symmetry. c.the graphs of two of these functions do not cross the x-axis d. the graphs of all these functions have different y- intercepts.

OpenStudy (eliesaab):

You can see that all the functions satisfy f(-x)=f(x) g(-x)=g(x) h(-x)=h(x) What does this tell you?

OpenStudy (mathmale):

Graphing all 3 of these functions would likely help you greatly. You'll need the graphing skills later as well.

OpenStudy (eliesaab):

Notice also that h(x) >=1, knowing this can the graph of h cross the x-axis

OpenStudy (eliesaab):

Which other function cannot cross the x-axis?

OpenStudy (eliesaab):

Notice that f(0)=6 g(0)=-5 h(0)=1 What does this tell you about their y-intercept

OpenStudy (eliesaab):

You should be able to finish the problem

OpenStudy (gudetamas):

they all have different y intercepts

OpenStudy (gudetamas):

@eliesaab

OpenStudy (mathmale):

What specifically do you need help with at this point?

OpenStudy (gudetamas):

so thats the answer correct?

OpenStudy (gudetamas):

@mathmale

OpenStudy (eliesaab):

By my remarks above, you should be able to see that b,c and d are correct